OpenStudy (anonymous):

what is the least common denominator of 2/x+x/x+1+x+1/3x

OpenStudy (anonymous):

\[\frac{ 2 }{ x }+\frac{ x }{ x+1 }+\frac{ x+1 }{ 3x }\]

OpenStudy (anonymous):

\[3x(x+1)\] is the least common multiple of the denominators